OpenStudy (anonymous):

The numbers x,y are chosen from {1,2,3........5N}.The probability that x^4-y^4 is divisible by 5 is?

OpenStudy (anonymous):

Find the generalised result i.e for all possible N

OpenStudy (anonymous):

\[x,y \in \mathbb{N} \:\: ?\]

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

This is tricky, and considering the fact I suck at probability, it gets even more tricky for me. \[x^4-y^4 = (x-y)(x+y)(x^2+y^2) \] Now I need to get solutions for each factor.

OpenStudy (anonymous):

How to do that?

OpenStudy (anonymous):

\[x^4 + y^4 = 5k \implies (x-y)(x+y)(x^2+y^2)=5k\] I'll ping @FoolForMath and @Mr.Math , maybe they could do it... I will need to think about it, it will take time for me.

OpenStudy (dominusscholae):

From what I can think, this problem says this: what is the probability that either x-y, x+y, or x^2+y^2 is a multiple of 5. for x^2+y^2, 3^2+4^2=5^2, so this leads me to believe that the two numbers picked in this case make this part of problem "what is the probability that two numbers picked in N with one number being a multiple of 3 and the other a multiple of 4. The only problem here are the other two cases...

OpenStudy (nikvist):

Can x^4-y^4 be negative?

OpenStudy (anonymous):

Yeah, I think so...

OpenStudy (nikvist):

\[17/25=0,68\]

OpenStudy (mr.math):

That's very smart nikvist! You're the hidden great mind in this site :)

OpenStudy (anonymous):

Solution comes as a function of N

OpenStudy (nikvist):

I think, solution does not depend on N.